OpenStudy (anonymous):

solve for the variable. Show work please. 4(x-2)=3^2-x

mathslover (mathslover):

first try to expand :4(x-2)

OpenStudy (anonymous):

4x-8

mathslover (mathslover):

great now we have : 4x -8 = \(\large{3^2-x}\) now what is 3^2 ?

OpenStudy (anonymous):

9

mathslover (mathslover):

right so we have : 4x - 8 = 9 - x right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

4x- 8 = 9-x add 8 both sides 4x - 8 + 8 = 9 - x + 8 4x = 9+8-x

mathslover (mathslover):

so 4x = 17 - x right?

OpenStudy (anonymous):

4x=17-x yes

mathslover (mathslover):

good add x both sides what do you get?

OpenStudy (anonymous):

5x=17

mathslover (mathslover):

good now divide 5 both sides

OpenStudy (anonymous):

3.4=x

mathslover (mathslover):

x = 17 / 5 = 3.4 very good

mathslover (mathslover):

So finally you got it?

OpenStudy (anonymous):

yes

mathslover (mathslover):

great to hear.

mathslover (mathslover):

also : \[\huge{\mathcal{Welcome}\space \mathcal{To} \space \mathcal{Openstudy}}\]